# Supersaturation of Tree Posets

### v1

@@ 3 Notation
We introduce some notation that will be crucial to our arguments.
Throughout our paper, we drop floors and ceiling whenever these are not crucial to our analysis. Using Chernoff bounds, it is standard [5, 16] to show that the number of sets $F\in\mathcal{B}_{n}$ with $||F|-n/2|>\sqrt{n\ln{n}}$ is $o({n\choose n/2})$. Thus, whenever we are given a dense subfamily $\mathcal{F}$ of $\mathcal{B}_{n}$, by leaving out at most $o(\binom{n}{n/2})$ members of $\mathcal{F}$, we may assume that all $F\in\mathcal{F}$ lie in the family
~$\widetilde{\mathcal{B}}_{n}:=\left\{F\in\mathcal{B}_{n}:\left|F-\frac{n}{2}\right|<\sqrt{n\ln n}\right\}.$
We note for later that any $F\in\widetilde{\mathcal{B}}_{n}$ satisfies
(${n\choose|F|}=\Omega(n^{-2}{n\choose n/2}),$ (1)
see for example [26, Equation 5.43].
Given a subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ and a positive integer $q$, we say that a tuple $(F_{1},\ldots,F_{q})$ of members of $\mathcal{F}$ is a $q$-chain if $F_{1}\supset F_{2}\supset\cdots\supset F_{q}$. We will abuse notation slightly by occasionally identifying decreasing tuples $(F_{1},\ldots,F_{q})$ by the corresponding set $\{F_{1},\ldots,F_{q}\}$. We use $\mathcal{C}$ to denote the family of all full chains of $\mathcal{B}_{n}$ (with $n$ fixed), i.e. $\mathcal{C}$ is the set of chains of length $n+1$ in $\mathcal{B}_{n}$.
Given a $q$-chain $Q$ in $\mathcal{F}$ and a full chain $\chi$ in $\mathcal{B}_{n}$ that contains all of the members of $Q$, we call the pair $(\chi,Q)$ a $q$-marked chain with markers in $\mathcal{F}$ or a $q$-marked chain from $\mathcal{F}$. Our proof will rely on obtaining a nice nested sequence of $q$-marked chains from $\mathcal{F}$ that are of a specific form. To this end, given a family $\mathcal{T}$ of $1$-marked chains of some subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ and $\chi\in\mathcal{C}$, we define $\mathcal{T}(\chi)=\{F:(\chi,F)\in\mathcal{T}\}$. We say that $\mathcal{T}$ is $q$-strong if for each $\chi\in\mathcal{C}$ where $\mathcal{T}(\chi)\neq\emptyset$, we have that $|\mathcal{T}(\chi)|\geq q$. For a $q$-strong $1$-marked chain family $\mathcal{T}$ from $\mathcal{F}$, we define the $q$-th power of $\mathcal{T}$, denoted by $\mathcal{T}[q]$, to be the $q$-marked chain family
@@
The nested families of $q$-marked chains that we build will be $q$-th power of some $q$-strong $1$-marked chain families.
Following standard notation in the literature, we define the Lubell weight $\mu(\mathcal{F})$ of a subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ to be
@@
4 Counting Lemmas on Marked Chains
@@
In this section, we collect some basic counting lemmas about marked chains that we will use in the final step of our proof Theorem 6.4. We begin by recalling the following lemma for $q$-marked chains.
Lemma 4.1 (Bukh [6], Lemma 4).
@@
Clearly, $\sum_{i<q}iC_{i}\leq(q-1)n!$. Hence, $|\mathcal{T}|=\sum_{i\geq q}iC_{i}\geq\varepsilon n!$. For the second statement, suppose $|\mathcal{F}|\geq(q-1+\varepsilon)\max_{F\in\mathcal{F}}\binom{n}{|F|}$. Then
@@
For this we recall that we order the members of our $q$-chain $(F_{1},\ldots,F_{q})$ by $F_{1}\supset F_{2}\supset\cdots\supset F_{q}$.
@@
5 Main Cleaning Result
As before, let $\mathcal{C}$ be the collection of all full chains $\chi$ in $\mathcal{B}_{n}$. Given a family $\mathcal{F}\subseteq\mathcal{B}_{n}$ and a family $\mathcal{M}$ of $q$-marked chains from $\mathcal{F}$, we define for each $i\in[q]$, $F\in\mathcal{F}$ and $\chi\in\mathcal{C}$ the sets
@@ 3. $(c)$
Any such subfamily $\mathcal{W}$ will be called an $(i,\delta)$-lower witness for $F$. In other words, $F$ being $(i,\delta)$-lower bad means that there exists a subfamily of small measure (in the sense of property (c).) $\mathcal{W}$ of members below $F$ such that every chain in $\mathcal{M}$ which has $F$ as the $i$th member must pass through this subfamily. While there may be many subfamilies $\mathcal{W}$ which are $(i,\delta)$-lower witnesses for $F$, in some contexts it will be useful to work with some fixed canonical witness. To this end, for any $F$ that is $(i,\delta)$-lower bad with respect to $\mathcal{M}$, we let $\mathcal{W}(F,i,\mathcal{M})$ denote the lexicographically minimal $\mathcal{W}$ which is an $(i,\delta)$-lower witness.
Note that if $\mathcal{W}$ is an $(i,\delta)$-lower witness, then property (a) implies
@@ 3. $(\hat{c})$
The rest of the section is dedicated to proving the following result which builds a nested sequence of families of $q$-marked chains with some robustness features with the additional property that each of these families is the $q$-th power of some family of $1$-marked chains (recall the definition of the $q$-th power of a $1$-marked chain family from (2).) This theorem provides the most important ingredient of our main proof.
@@
For all integers $q\geq 1$ and for all reals $\varepsilon>0$, there exists some $\delta>0$ such that the following holds. Let ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ and let $\mathcal{T}^{0}$ be a $q$-strong $1$-marked chain family with markers from $\mathcal{F}$ such that $|\mathcal{T}^{0}|\geq\varepsilon n!$. Then there exists a collection of $\mathcal{T}^{|P|}\subseteq\mathcal{T}^{|P|-1}\subseteq\dots\subseteq\mathcal{T}^{0}$
@@
Before starting the proof, we set up some notation and claims that are useful in the main proof. In what follows, we fix ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ as in the theorem statement together with the large constant
@@
Starting with $\mathcal{T}^{0}$, we will build our subsets $\mathcal{T}^{j}$ as follows. Suppose $\mathcal{T}^{j-1}$ has already been defined for some $1\leq j\leq|P|$, and for ease of notation let $\mathcal{M}^{j-1}=\mathcal{T}^{j-1}[q]$. For each $i=1,\dots,q$ and for every $\chi\in\mathcal{C}$, let $B^{j-1}(\chi,i,\uparrow)$ be the set of members $F$ in $\mathcal{T}^{j-1}(\chi)$ such that $F$ is the $i$th member of some $q$-chain in $\mathcal{M}^{j-1}(\chi)$ and $F$ is $(i,\delta)$-upper bad to $\mathcal{M}^{j-1}$. Let $B^{j-1}(\chi,i,\downarrow)$ be the set of members $F$ in $\mathcal{T}^{j-1}(\chi)$ such that $F$ is $i$th member of some $q$ chain in $\mathcal{M}^{j-1}(\chi)$ and $F$ is $(i,\delta)$-lower bad with respect to $\mathcal{M}^{j-1}$. For convenience we denote the union of these sets by
@@
$\mathcal{T}^{j}=\{(\chi,F):\chi\in\mathcal{C}_{2}^{j},\ F\in\mathcal{T}^{j-1}(\chi)-B^{j-1}(\chi)\}$
@@
$\operatorname{Prob}[A_{1}\cap A_{2}\cap\dots\cap A_{m}]$ $=\operatorname{Prob}[A_{1}]\operatorname{Prob}[A_{2}|A_{1}]\dots\operatorname{Prob}[A_{m}|A_{1}\cap A_{2}\cap\dots\cap A_{m-1}]$
@@
since conditioned on $\chi_{\mathcal{F}}(c_{\ell})=F$ the events $\chi_{\mathcal{F}}(c_{i}^{\prime})\in\mathcal{W}^{*}(F)$ and $A_{1}\cap A_{2}\cap\dots A_{\ell-1}$ are independent. Thus, by definition of $\mathcal{W}^{*}(F)$, we have
@@
Observe that $A_{\ell}$ is exactly the union over the events $\chi_{\mathcal{F}}(c_{\ell})=F$ over $F\in\mathcal{S}$ and $\chi_{\mathcal{F}}(c_{\ell}^{\prime})\in\mathcal{W}^{*}(F)$. Since these events are disjoint, we have that the following inequalities holds
@@
$\leq\sum_{F\in\mathcal{S}}\operatorname{Prob}[\chi_{\mathcal{F}}(b_{\ell})=F|A_{1}\cap A_{2}\cap\dots A_{\ell-1}]\delta$
$\leq\delta$
@@
Finally, we combine all of the results we have established up to this point into a single statement. For this we recall that $\mu(\mathcal{F})=\sum_{F\in\mathcal{F}}\frac{1}{{n\choose|F|}}$ denotes the Lubell weight of $\mathcal{F}$, that
$\mathcal{L}^{i}(\mathcal{M})=\{D\in\mathcal{B}_{n}:D\text{ is the }i\text{th member of a }q\text{-chain in }\mathcal{M}\},$
and the definition of $\delta$-robustness defined at the start of this section.
@@
6 Proof of Theorem 1.8
Throughout this section, $P$ be a fixed tree poset of height $k\leq q$. Given a set $F\in\mathcal{B}_{n}$, we let $U(F)=\{S\in\mathcal{B}_{n}:S\supseteq F\}$ and $D(F)=\{S\in\mathcal{B}_{n}:S\subseteq F\}$. If $\mathcal{S}$ is a subfamily of $B_{n}$ we let

### v2

@@ 2 Notation
Throughout our paper, we drop floors and ceiling whenever these are not crucial to our analysis. Let
~$\widetilde{\mathcal{B}}_{n}:=\left\{F\in\mathcal{B}_{n}:\left|F-\frac{n}{2}\right|<2\sqrt{n\ln n}\right\}.$
Using some standard tools, we will assume that the family $\mathcal{F}$ we are working with is subfamily of ~$\widetilde{\mathcal{B}}_{n}$ instead of $\mathcal{B}_{n}$.
We define the Hasse diagram of $P$, denoted $H(P)$, as a directed graph with vertex set $P$ where there is an edge from $x$ to $y$ only if $y>x$ and there is no $z$ such that $y>z>x$. Notice that this definition differs slightly from the classical definition of Hasse diagram via undirected graphs embedded on the plane, but for us it will be more convenient to use the directed setting.
Given a subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ and a positive integer $q$, we say that a tuple $(F_{1},\ldots,F_{q})$ of members of $\mathcal{F}$ is a $q$-chain if $F_{1}\supset F_{2}\supset\cdots\supset F_{q}$. In particular, when we refer to the $i$th element of any $q$-chain we mean the one with $i$th largest cardinality. We will abuse notation slightly by occasionally identifying decreasing tuples $(F_{1},\ldots,F_{q})$ by the corresponding set $\{F_{1},\ldots,F_{q}\}$. We use $\mathcal{C}$ to denote the family of all full chains of $\mathcal{B}_{n}$ (with $n$ fixed), i.e. $\mathcal{C}$ is the set of chains of length $n+1$ in $\mathcal{B}_{n}$. The height of $P$ is the largest length of a chain in $P$.
Given a $q$-chain $Q$ in $\mathcal{F}$ and a full chain $\chi$ in $\mathcal{B}_{n}$ that contains all of the members of $Q$, we call the pair $(\chi,Q)$ a $q$-marked chain with markers in $\mathcal{F}$ or a $q$-marked chain from $\mathcal{F}$. Given a family of $q$-marked chains $\mathcal{M}$, let
$\mathcal{L}^{i}(\mathcal{M})=\{D\in\mathcal{B}_{n}:D\text{ is the }i\text{th member of a }q\text{-chain in }\mathcal{M}\}.$
Our proof will rely on obtaining a nice nested sequence of $q$-marked chains from $\mathcal{F}$ that are of a specific form. To this end, given a family $\mathcal{T}$ of $1$-marked chains of some subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ and $\chi\in\mathcal{C}$, we define $\mathcal{T}(\chi)=\{F:(\chi,F)\in\mathcal{T}\}$. We say that $\mathcal{T}$ is $q$-strong if for each $\chi\in\mathcal{C}$ where $\mathcal{T}(\chi)\neq\emptyset$, we have that $|\mathcal{T}(\chi)|\geq q$. For a $q$-strong $1$-marked chain family $\mathcal{T}$ from $\mathcal{F}$, we define the $q$-th power of $\mathcal{T}$, denoted by $\mathcal{T}[q]$, to be the $q$-marked chain family
@@
We will consider some of our results in a somewhat more general setting. To this end, we define the Lubell weight of a family $\mathcal{F}$, denoted by $\mu(\mathcal{F})$, by
@@
The Lubell weight is a natural measure to put on families in $\mathcal{B}_{n}$, and in particular we note that $\mu(\mathcal{F})$ is the expected number of members of $\mathcal{F}$ that are contained in a uniformly randomly chosen full chain. While this notion is not strictly needed for our work, we are able to apply our methodology in this more general setting which has historically been of interest to extremal problems for posets, see for example [29].
In Section 6, we will use the Chernoff’s Inequality in the following form, see [22]:
Lemma 2.1.
If $X=\mathrm{Bin}(n,p)$ and $\delta\leq 1$,
$\operatorname{Prob}(|X-\mathbb{E}[X])|\geq\delta\mathbb{E}[X]\leq 2\exp\left(-\frac{\delta^{2}}{3}\mathbb{E}[X]\right).$
3 Sketch of The Proof of Theorem 1.3
Here we sketch how to prove Theorem 1.3 which we recall says that every large family $\mathcal{F}\subset\mathcal{B}_{n}$ contains many copies of a given tree poset $P$. We will do this by showing that there are many ways of embedding $P$ into $\mathcal{F}$, but before we get into this we need some preliminaries.
It will be convenient to associate each member $F$ of a given subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ with the full chains $\chi$ of $\mathcal{B}_{n}$ that contain it, so we will work with pairs $(\chi,F)$ instead of individual members $F$, where $F\in\mathcal{F}$ and $\chi$ is any full chain in the boolean lattice. Likewise, we will associate each $q$-chain $Q$ in our family $\mathcal{F}$ with the full chains $\chi$ of $\mathcal{B}_{n}$ that contain it and call such pairs $(\chi,Q)$ $q$-marked chains. It is worth mentioning that the approach of working with $q$-marked chains was originated by Bukh [9] and was later further developed by Boehnlein and Jiang [8].
Using Chernoff bounds, it is standard [8, 21] to show that the number of sets $F\in\mathcal{B}_{n}$ with $||F|-n/2|>2\sqrt{n\ln{n}}$ is $o({n\choose n/2})$. Thus, whenever we are given a dense subfamily $\mathcal{F}$ of $\mathcal{B}_{n}$ i.e. $|\mathcal{F}|=\Omega({n\choose n/2})$, by leaving out at most $o(\binom{n}{n/2})$ members of $\mathcal{F}$, we may assume that all $F\in\mathcal{F}$ lie in the family ~$\widetilde{\mathcal{B}}_{n}$.
At the heart of all of our arguments is a general result Theorem 4.15 which says that any subset ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ of large Lubell weight, more precisely at least of weight at least $q-1+\varepsilon$, contains a nested sequence of large $q$-marked chain families that have a certain robustness property which we will describe momentarily. Note that if $\mathcal{F}$ is of size at least $(q-1+\varepsilon){n\choose\lfloor n/2\rfloor}$ then the hypothesis above on the Lubell weight is satisfied.
Before the embedding starts, we fix an ordering $x_{1},x_{2},\ldots,x_{|P|}$ of vertices of $P$ such that each $x_{j+1}$ has a unique neighbour (called the parent) among $x_{1},\dots x_{j}$, and we fix a poset homomorphism $\mathrm{r}:P\rightarrow[q]$, with $[q]$ under the reverse of the natural total ordering. As a pre-processing step, we use our main tool Theorem 4.15 to generate a nested family of $q$-marked chains $\mathcal{M}^{0}\supseteq\mathcal{M}^{1}\supseteq\cdots\supseteq\mathcal{M}^{|P|}$ such that if $F\in\mathcal{F}$ is in some $q$-marked chain of $\mathcal{M}^{i}$ family then it is “robust” with respect to $\mathcal{M}^{i-1}$ as well, meaning $F$ is contained in “many” $q$-chains of $\mathcal{M}^{i-1}$ as well. The idea of nested families $\mathcal{M}^{i}$ may seem a bit peculiar but unfortunately we were not able to guarantee the existence of a single marked chain family $\mathcal{M}$ with the property of every $F$ being contained in the desired number of copies of $q$-marked chains in $\mathcal{M}$. To obtain this nested sequence, we iteratively remove all the sets $F\in\mathcal{F}$ which are bad with respect to $\mathcal{M}^{j-1}$ and the relevant $q$-chains from $\mathcal{M}^{j-1}$ subsequently, and ensure that the sizes of $\mathcal{M}^{j}$ do not shrink dramatically, thus guaranteeing that after $|P|$ many steps we still have a large family of $q$-marked chains.
Having run the cleaning process, we embed the vertices $x_{1},x_{2},\dots x_{|P|}$ iteratively so that at the $j$th step we find many partial embedding of the form $\varphi^{j}:\{v_{1},v_{2},\dots,v_{j}\}$. We embed $v_{1}$ arbitrarily in the $\mathrm{r}(v_{1})$-level of any $q$-marked chain inside $\mathcal{M}^{|P|}$. Notice that the number of elements in the $\mathrm{r}(v_{1})$-level of a chain in $\mathcal{M}^{|P|}$ gives us a lower bound on the number of ways to embed $v_{1}$. Also while $\mathcal{M}^{|P|}$ is the smallest family of $q$-marked chains it has the property that every $F$ in a $q$-chain of $\mathcal{M}^{|P|}$ is “robust” which respect to every $\mathcal{M}^{j}$ with $j<|P|$. This is the incentive behind embedding the vertices of $P$ in the reverse order of the nested families, more precisely, $v_{j}$ will be embedded in $\mathcal{M}^{|P|-j+1}$. At the $j$th step, we embed $v_{j}$, assuming that its parent $y$ who is among $\{v_{1},v_{2},\dots,v_{j-1}\}$ is already embedded. We wish to embed $v_{j}$ in the $\mathrm{r}(v_{j})$th position of some $q$-marked chain of $\mathcal{M}^{|P|-j+1}$ which has $\varphi^{j-1}(y)$ in the $\mathrm{r}(y)$th position. By the cleaning process we ran earlier, we are guaranteed to have many such choices to embed $v_{j}$ and find many partial embeddings $\varphi^{j}$. This embedding procedure in total gives a lower bound on the number of induced copies of $P$ in $\mathcal{F}$ as a function of our choice of $\mathrm{r}$. It turns out there is a relation between $M^{*}(n,q,P)$ and the counting of copies of $P$ via all such rank functions $\mathrm{r}$ (see Lemma 5.5). Thus we may choose $\mathrm{r}$ suitably and get this lower bound to be a constant fraction of $M^{*}(n,q,P)$. This gives us our desired supersaturation result saying that the number of induced copies of $P$ guaranteed in $\mathcal{F}$ is as least as large as a fraction of the number of induced copies of $P$ in the middle $q$-levels without knowing $M^{*}(n,q,P)$ explicitly.
The full details of this embedding are slightly technical, as one also needs to make sure that the embedding guarantees noncomparable pairs stay noncomparable and other nuances, however the main gist of the argument is that the cleaning process in Theorem 4.15 provides the framework to do this successfully.
4 Key Tools
In this section, we state and prove our main tool Theorem 4.15, which roughly speaking guarantees in any large $\mathcal{F}\subseteq\mathcal{B}_{n}$ the existence of a nested sequence of $q$-marked chains $\mathcal{M}^{|P|}\subseteq\mathcal{M}^{|P|-1}\subseteq\cdots\subseteq\mathcal{M}^{0}$ which are “robust” in a certain way. We begin by establishing some preliminary lemmas.
4.1 Counting Lemmas on Marked Chains
@@
In this section, we collect some basic counting lemmas about marked chains that we will use in the final step of our proof Theorem 5.4. We begin by recalling the following lemma of Bukh [9, Lemma 4] for $q$-marked chains.
Lemma 4.1 ([9]).
@@
Clearly, $\sum_{i<q}iC_{i}\leq(q-1)n!$. Hence, $|\mathcal{T}|=\sum_{i\geq q}iC_{i}\geq\varepsilon n!$.
For the second statement, suppose $|\mathcal{F}|\geq(q-1+\varepsilon)\max_{F\in\mathcal{F}}\binom{n}{|F|}$. Then
@@
4.2 Main Cleaning Result
Given a family $\mathcal{F}\subseteq\mathcal{B}_{n}$ and a family $\mathcal{M}$ of $q$-marked chains from $\mathcal{F}$, we define for each $i\in[q]$, $F\in\mathcal{F}$ and $\chi\in\mathcal{C}$ the sets
@@ 3. $(c)$
Any such subfamily $\mathcal{W}$ will be called an $(i,\delta)$-lower witness for $F$.
Informally, $F$ being $(i,\delta)$-lower bad means that there exists a subfamily of small measure (in the sense of property (c)) $\mathcal{W}$ of members below $F$ such that every chain in $\mathcal{M}$ which has $F$ as the $i$th member must pass through this subfamily. While there may be many subfamilies $\mathcal{W}$ which are $(i,\delta)$-lower witnesses for $F$, in some contexts it will be useful to work with some fixed canonical witness. To this end, for any $F$ that is $(i,\delta)$-lower bad with respect to $\mathcal{M}$, we let $\mathcal{W}(F,i,\mathcal{M})$ denote the lexicographically minimal $\mathcal{W}$ which is an $(i,\delta)$-lower witness. Note that for all $\mathcal{W}$,
@@ 3. $(\hat{c})$
The rest of the section is dedicated to proving the following result which builds a nested sequence of families of $q$-marked chains with some robustness features with the additional property that each of these families is the $q$-th power of some family of $1$-marked chains (recall the definition of the $q$-th power of a $1$-marked chain family from (1)). This theorem provides the most important ingredient of our proof of Theorem 4.15.
@@
For all integers $q\geq 1$ and for all reals $\varepsilon>0$, there exists some $\delta>0$ such that the following holds. Let ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ and let $\mathcal{T}^{0}$ be a $q$-strong $1$-marked chain family with markers from $\mathcal{F}$ such that $|\mathcal{T}^{0}|\geq\varepsilon n!$. Then there exists a collection of $1$-marked chains $\mathcal{T}^{|P|}\subseteq\mathcal{T}^{|P|-1}\subseteq\dots\subseteq\mathcal{T}^{0}$ satisfying the following:
@@
Before starting the proof, we set up some notation and claims that will be useful. In what follows, we fix ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ as in the theorem statement together with the large constant
@@
Starting with $\mathcal{T}^{0}$, we will build our subsets $\mathcal{T}^{j}$ as follows. Suppose $\mathcal{T}^{j-1}$ has already been defined for some $1\leq j\leq|P|$, and for ease of notation let $\mathcal{M}^{j-1}=\mathcal{T}^{j-1}[q]$. For each $i=1,\dots,q$ and for every $\chi\in\mathcal{C}$, let $B^{j-1}(\chi,i,\uparrow)$ be the set of members $F$ in $\mathcal{T}^{j-1}(\chi)$ such that $F$ is the $i$th member of some $q$-chain in $\mathcal{M}^{j-1}(\chi)$ and $F$ is $(i,\delta)$-upper bad to $\mathcal{M}^{j-1}$. Let $B^{j-1}(\chi,i,\downarrow)$ be the set of members $F$ in $\mathcal{T}^{j-1}(\chi)$ such that $F$ is $i$th member of some $q$-chain in $\mathcal{M}^{j-1}(\chi)$ and $F$ is $(i,\delta)$-lower bad with respect to $\mathcal{M}^{j-1}$. For convenience we denote the union of these sets by
@@
$\mathcal{T}^{j}=\{(\chi,F):\chi\in\mathcal{C}_{2}^{j},\ F\in\mathcal{T}^{j-1}(\chi)-B^{j-1}(\chi)\}.$
@@
$\operatorname{Prob}[A_{1}\cap A_{2}\cap\dots\cap A_{m}]$ $=\operatorname{Prob}[A_{1}]\operatorname{Prob}[A_{2}|A_{1}]\dots\operatorname{Prob}[A_{m}|A_{1}\cap A_{2}\cap\dots\cap A_{m-1}].$
@@
since conditioned on $\chi_{\mathcal{F}}(c_{\ell})=F$ the events $\chi_{\mathcal{F}}(c_{i}^{\prime})\in\mathcal{W}^{*}(F)$ and $A_{1}\cap A_{2}\cap\dots A_{\ell-1}$ are independent.
Thus, by definition of $\mathcal{W}^{*}(F)$, we have
@@
Observe that $A_{\ell}$ is exactly the union over the events $\chi_{\mathcal{F}}(c_{\ell})=F$ over $F\in\mathcal{S}$ and $\chi_{\mathcal{F}}(c_{\ell}^{\prime})\in\mathcal{W}^{*}(F)$. Since these events are disjoint, we have that the following series of inequalities hold:
@@
$\leq\sum_{F\in\mathcal{S}}\operatorname{Prob}[\chi_{\mathcal{F}}(c_{\ell})=F|A_{1}\cap A_{2}\cap\dots A_{\ell-1}]\delta$
$\leq\delta.$
@@
Finally, we combine all of the results we have established up to this point into a single statement. Also note that the second part of this result is not needed for our proofs, but we include it in the statement since it adds no extra difficulty to the proof and may be useful for future applications.
@@
4.3 Tools for Induced Posets
Throughout this subsection, let $P$ be a fixed tree poset of height $k\leq q$. Given a set $F\in\mathcal{B}_{n}$, we let $U(F)=\{S\in\mathcal{B}_{n}:S\supseteq F\}$ and $D(F)=\{S\in\mathcal{B}_{n}:S\subseteq F\}$. If $\mathcal{S}$ is a subfamily of $B_{n}$ we let